OpenStudy (anonymous):

integral of (x-1)e^x/x^2?

OpenStudy (anonymous):

1) distribute 2) split fraction into 2 integrals 3) solve first integral 4) use by parts on second integral with dv=1/x^2 5) use by parts again with dv=1/x

OpenStudy (anonymous):

integral of e^x/x?

OpenStudy (anonymous):

yup... by parts again... with dv=1/x

OpenStudy (anonymous):

no but for the first integral, not the second. Do I use parts 3 times?

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{1}{x}dx-\int\limits_{}^{}\frac{e ^{x}}{x ^{2}}dx\]

OpenStudy (anonymous):

isn't the first one supposed to be integral of e^x/x, not 1/x?

OpenStudy (anonymous):

\[\ln \left| x \right|+\frac{2e ^{x}}{x}+\int\limits_{}^{}\frac{e ^{x}}{x}dx\]

OpenStudy (anonymous):

are you looking at\[\int\limits_{}^{}(x-1)\frac{e ^{x}}{x ^{2}}dx\]?

OpenStudy (anonymous):

oops... yes... by parts on the first one too then

OpenStudy (anonymous):

yes, which is integral of xe^x/x^2 - integral of e^x/x^2

OpenStudy (anonymous):

u = e^x or does u =1/x?

OpenStudy (anonymous):

u=e^x

OpenStudy (anonymous):

that's what i just did and it looks like im in a loop.

OpenStudy (anonymous):

yuck... I agree

OpenStudy (anonymous):

ill try making 1/x = u

OpenStudy (anonymous):

If you do the other by parts twice... I think that you might get something from it though.

OpenStudy (anonymous):

Sorry.. I'm thinking backward: \[\frac{d}{dx}(\frac{e^x}{x}) = \frac{xe^x - e^x}{x^2}\]So....

OpenStudy (anonymous):

another loop???

OpenStudy (anonymous):

\[\int\limits_{ }^{ } e ^{x} *\frac{ 1 }{ x} = \frac{ e ^{x} }{ x } -\int\limits_{ }^{ }-e ^{x} *\frac{ 1 }{ x ^{2} }\]

OpenStudy (anonymous):

add integral e^x * -1/x^2 to both sides

OpenStudy (anonymous):

yes... I am getting one too.... working.

OpenStudy (anonymous):

there we go.. it's recursive

OpenStudy (anonymous):

got that, but when you do parts again on that answer algebraic! i get another loop since the new integral is integral of vdu = 2*integeral of e^x/x^3

OpenStudy (anonymous):

naw you're done.

OpenStudy (anonymous):

think about it for a bit.

OpenStudy (anonymous):

PM me when 'eureka'

OpenStudy (anonymous):

oh wait so its just e^x/x?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

haha wow i kept trying to integrate but the integrals are opposites of each other. Thank you so much.

OpenStudy (anonymous):

glad to be of modest service:)

OpenStudy (anonymous):

awww man.... nice call